1) Prove following Fourier transform: x(t)cos (Wot+0) 3 jx(w-wo)el® + X(w+w)e=;8]
Answer #1

KURA Given xH) cos(wot to) steps = xH) / coswot coso - sin wot sino ) (Formula : cos(A+B) = COSA COSB -sina sinB) steps = xltFormula: stereo xit) cio Cicerot) tera lefoot)) Applying fourier transform swot = 8 (w.wo); e wot = 8 (wtwo), = X(W) / ( scw-